OpenStudy (anonymous):

write an equation of the line containing the given point and perpendicular to the given line. express your answer in the form y=mx+b (-2,6);6x=5y+8

OpenStudy (anonymous):

The given line has slope 6/5 so the slope of the perpendicular has slope -5/6. Then, y = -5/6(x-(-2)) + 6 is a line with that slope containing (-2,6) y = -5x/6 - 10/6 + 6 = (-5x + 26)/6